OpenStudy (anonymous):

Ms. Yamato’s gross pay is $2644. Her deductions total $548.30. What percent of her gross pay is take-home pay?

OpenStudy (anonymous):

To find her take home pay, you subtract her deductions from her gross pay and then divide her take-home pay by her gross pay and then multiply by 100%.

OpenStudy (anonymous):

Take-home pay = 2644-548.30=2095.70 Percent = (2095.70/2644)*100%=79.262%

OpenStudy (anonymous):

According to this partial W-2, how much money was paid in FICA taxes?